# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v4

@@ Abstract
Fiducial markers are widely used in various robotics tasks, facilitating enhanced navigation, object recognition, and scene understanding. While offering significant advantages for robots and Augmented Reality (AR) applications, they normally disrupt the visual aesthetics of environments because they are visible to humans, making them unsuitable for many everyday use cases. To address this gap, this paper presents iMarkers—innovative, unobtrusive fiducial markers detectable exclusively by robots and AR devices equipped with adequate sensors and detection algorithms. These markers offer high flexibility in production, allowing customization of their visibility range and encoding algorithms to suit various demands. The paper also introduces the hardware designs and open-sourced software algorithms developed for detecting iMarkers, highlighting their adaptability and robustness in the detection and recognition stages. Various evaluations have demonstrated the effectiveness of iMarkers compared to conventional (printed) and blended fiducial markers and confirmed their applicability in diverse robotics scenarios.
@@ 1 Introduction
Fiducial markers, i.e., artificial landmarks with distinguishable graphical patterns, are widely used in robotics and computer vision applications to offer reliable object identification, localization, and pose estimation [1]. As markers provide better-defined reference points than naturally available ones, they are simple-yet-effective solutions where consistent feature-matching and precise pose retrieval are essential. Additionally, augmenting information via markers placed in environments and employing proper systems to extract that information can benefit a wide range of cases, including Augmented Reality (AR) and robotics tasks [2]. Markers can narrow the range of objects a robot needs to recognize and provide real-time pose information for tracking and localization. For instance, decoding data from markers detected using the robot’s camera is less costly than processing the entire visual scene that the robot can see.
Figure 1: A legged robot observing an environment labeled with several printed fiducial markers and the proposed iMarkers: a) the original scene, b) the same scene with iMarkers highlighted and magnified.
Although such markers can convey various advantages in robotics, seeing multiple objects labeled with printed markers can be aesthetically undesirable and destructive. In other words, many would want to avoid seeing printed tags around them and prefer that the robots operate well without spoiling zones with these artificial landmarks. Moreover, evidence has shown that visible markers can disrupt and interrupt natural gaze behaviors [3]. Consequently, the authors worked on the idea of employing Cholesteric Spherical Reflectors developed by the team to produce a new generation of fiducial markers [4], dubbed iMarker [^1] https://snt-arg.github.io/iMarkers/. This can be done by replacing the ink pigment used in printing common fiducial markers with CSRs of spherical shell type. Due to the peculiar behavior of CSRs, iMarkers can be invisible (indistinguishable) to human eyes but observable by robots in case a dedicated sensor is provided [5]. Importantly, the sensor can be made at a very low cost. Moreover, such markers can be fabricated to be detected only within specific wavelength ranges of light, thereby avoiding visual clutter in the environment. The iMarker concept is illustrated in Fig. 1, confirming, through qualitative comparison, that they are hardly noticeable by the naked eye and in standard camera images compared to traditional printed fiducial markers.
The primary motivation of this paper is to demonstrate the potential of iMarkers in robotics by optimizing their design and detection procedure and benchmarking possible use cases in real-world robotic applications. Moreover, another objective is verifying the potential of such markers to supply simplified solutions for complex computer vision tasks in robotics, such as detecting transparent and reflective objects (e.g., glass, windows, and mirrors). Hence, the leading contributions of the paper are summarized below:
@@
  • Proposing multiple sensor solutions and software algorithms for detecting iMarkers of various designs and extracting their 6DoF pose information; and
  • Assessing their potential applicability and effectiveness compared to traditional fiducial markers.
@@ 2.1 Ordinary Fiducial Markers
Fiducial marker libraries used in robotics and AR applications are expected to enable rapid detection, reliable recognition, and efficient information decoding. Ordinary fiducial markers, mainly printed on paper, can be classified into non-square (designed as circles, point sets, or arbitrary visual patterns), square (or matrix-based), and hybrid variations. In non-square markers, identification and data decoding generally occur w.r.t. the center of their geometric shape. InterSense [6] and WhyCode [7] are circular-shaped libraries in which the correspondence point detection is the tricky part. Dot-shaped libraries, such as RUNE-Tag [8], supply fast detection but suffer primarily in noisy conditions. Other arbitrary and shape-free markers are mainly used in 6-Degrees-of-Freedom (6DoF) feature tracking and pose estimation. Despite robust detection, the mentioned method faces challenges like inter-marker encoded dot confusion and steep viewing angle recognition.
Square fiducial markers, on the other hand, employ binary codes encoded in matrices and deliver four principal corner points for more precise pose estimation. They require less computation cost for detection and recognition, leading to potentially faster detection algorithms. ARToolkit [9] is an open-source matrix-based marker library that uses template-matching for square marker pattern determination. However, it may face high re-projection errors with rotation and camera angle changes. AprilTag [10] is another marker library that incorporates gradient-based edge detection and reliable line identification to ensure accurate marker recognition, even under challenging viewing angles. Its enhanced digital coding system has made it more robust to rotation, occlusion, and warping compared to traditional systems, such as ARToolkit. Similarly, Garrido-Jurad et al. introduced ArUco Marker [11] that generates configurable marker dictionaries with maximized inter-marker distances. Their system contains automatic marker detection with error correction based on dictionary search and a novel occlusion handling method using color-based occlusion masks for robust pose estimation.
Despite their widespread use and robust performance, these markers, whether non-square or square, remain visually intrusive and can disrupt environments’ natural aesthetics. This limitation has motivated the development of markers that blend more seamlessly into their surroundings.
@@ 2.2 Blended and Unobtrusive Fiducial Markers
In contrast with ordinary paper-based markers, some solutions propose fiducial markers for particular scenarios that require special sensors to be detected. In this regard, ArTuga [12] is a multimodal marker that targets detection in challenging environments by photometric and radiometric sensor fusion, mainly designed to enhance the precise landing of drones in harsh environments. Seedmarkers [13] consider the aesthetic properties of fiducial markers by embedding them in physical objects, ranging from laser-cut plates to 3D-printed tangibles. ACMarker [14] is another solution to create fiducial markers detectable using customized sensors for underwater robotics applications.
Several approaches focus on integrating less obtrusive markers into environments to minimize their visual impact while maintaining robust detectability. Transparent Random Dot Marker [15] is an approach that aims to keep markers on transparent sheets with small randomly placed printed dots, making them barely visible in ordinary scenarios. Nevertheless, geometric point matching and marker detection can be tricky while integrating pairwise relationships among the detected dots. InfraredTags [16] are discreet and durable markers designed to be printed and embedded within various objects, offering invisibility and rapid scanning. However, their production is limited to the Infrared (IR) range, restricting flexibility in fabrication. Additionally, their thin and embedded nature makes them difficult for users to locate, complicating the process of directing IR torches or cameras for detection. Similarly, BrightMarkers [17] are 3D-printed fluorescent filaments embedded into objects, which require Near Infrared (NIR) cameras for detection. Aircode [18] introduced the concept of embedding markers directly into the fabrication of objects, enabling seamless integration for use cases such as robotic grasping. Nevertheless, Aircode’s significant drawback is its slow decoding process, which can take tens of seconds, depending on the camera’s viewing angle.
@@ 2.3 Gap Identification
Although the surveyed blended and unobtrusive methodologies have indicated effectiveness in robotics and AR tasks by keeping markers discreet, their applicability in real-world scenarios is often restricted by the trade-off between invisibility and detection speed. In other words, solutions like AirCode achieve high levels of invisibility but suffer from significantly slow detection speeds. Other crucial factors include simplicity, cost-effectiveness, detection versatility, and robust performance, as well as fabrication flexibility across various light spectra. These dimensions significantly influence the adoption of such fiducial marker solutions compared to the straightforward and easily printable traditional markers.
@@
iMarkers [5] present an innovative and versatile solution to address the mentioned gaps by ensuring transparency, flexibility, and unobtrusiveness. Given the proper sensor and detection algorithm, they can be fabricated to achieve high contrast for detection while remaining undetectable to the naked eye. Thus, they improve fiducial markers’ functionality and visual harmony in various settings. Table 1 highlights the characteristics of iMarkers compared to other methodologies aimed at seamlessly blending markers into their surroundings. According to the table, although all solutions aim to remain hidden from human perception, iMarkers cover distinct aspects that set them apart, as outlined below:
@@
Core Material: The primary material used to make iMarkers is CSRs, microscopic droplets of Cholesteric Liquid Crystals (CLCs) that self-organize into a helical structure, making them highly effective at reflecting specific wavelengths of light, including IR, Ultraviolet (UV), or visible [5]. Their spherical shape turns them into omnidirectional selective retroreflectors, ensuring they can be easily detected regardless of viewing angle. Thus, they can produce vivid, circularly polarized iridescent colors resulting from the selective reflection of light. As shown in Fig. 2, iMarkers combine CSRs for the marker patterns with transparent UV-curable glue to maintain the structure of the fiducial marker. An alternative fabrication method involves directly spraying CSR shells onto surfaces, a promising approach currently under investigation.
Figure 2: The role of CSRs in the design of iMarkers: (a) an iMarker featuring ArUco patterns filled with CSRs, (b) an iMarker displayed next to visible CSR particles, illustrating various shades of green for fabrication.
@@
Fabrication Efficiency: The cost and simplicity of fabrication significantly impact the widespread adoption of fiducial markers. While AirCode requires partitioning and InfraStructs depend on complex, layered assembly, iMarkers, as shown in Fig. 2, can be fabricated using cost-effective methods like dispensing or spray coating.
Production Versatility: In contrast with other solutions, iMarkers offer production flexibility, enabling design and detection across the entire light spectrum. UV-range iMarkers remain entirely transparent and invisible to the naked eye, even when positioned on a transparent surface like glass. IR-range iMarkers are similarly discreet, though slight scattering or subtle color effects may reveal their presence (to some extent) under extreme lighting conditions. Even in the visible range, iMarkers can be designed to blend with patterned backgrounds or camouflage against similarly colored surfaces, ensuring minimal visual intrusion. Fig. 3 depicts actual examples of iMarkers made from different CSR shells.
Sensor Affordability: Detection requirements differ among unobtrusive marker solutions: AirCode relies on a large projector and camera setup, InfraStructs require costly Terahertz scanners, and InfraredTags depend on specialized imaging modules with microprocessors. In contrast, iMarkers can be detected using affordable sensors with standard optical components. The camera for detecting iMarkers must be compatible with the polarization and wavelength characteristics of the CSRs employed in its fabrication.
@@ 4 Sensor Design and Detection Strategies
Considering the versatility of iMarkers, a range of perception sensors and detection algorithms can be employed to ensure reliable recognition across diverse applications. Hence, three primary solutions are proposed for iMarker detection, including dual-vision, dynamic single-vision, and static single-vision, which are further described in the following:
@@ 4.1 Dual-vision Solution
Figure 4: Dual-vision setup variants designed for iMarker detection, each containing two synchronized cameras with attached circular polarizers to their lenses facing a cube (a) or plate (b) beamsplitter.
Sensor Design. The first sensor is a homogeneous perception system containing two perpendicularly positioned synchronized cameras, each facing a different surface of an optical beamsplitter. As shown in Fig. 4, the sensor can be assembled in two variants: one using a cube and the other employing a plate beamsplitter. The beamsplitter divides the incoming light equally (i.e., $50\texttt{:}50$) into two parts, including the transmitted (that passes through) and the reflected light. Thus, by mounting the cameras and beamsplitter within a cage setup, one camera captures the scene using the transmitted light, while the other captures it using the reflected light, resulting in two perspectives with minimal variation. Attaching identical circular polarizers to each camera’s lens effectively blocks light with the opposite circular polarization, ensuring only the desired handedness is detected.
Due to the inversion of circular polarization upon reflection in the beamsplitter, the camera capturing the reflected light detects the “opposite” circular polarization compared to the camera capturing the transmitted light. Accordingly, the sensor simultaneously captures the same scene with two cameras, ensuring that the CSR regions of iMarkers in the sensor’s Field of View (FoV) appear blocked in one camera while remaining visible in the other (due to the opposing circular polarization of the CSR reflections).
The cameras in this setup must be synchronized to capture the same scene simultaneously. Simplistically, frame $f_{t}^{C_{1}}$ captured by camera $C_{1}$ at time $t$ should depict the same scene as frame $f_{t}^{C_{2}}$ of camera $C_{2}$. Additionally, as the cameras’ frame rate and shutter speed are vital factors for high-speed robotic applications (e.g., drone navigation), integrating global shutter cameras in this sensor is highly recommended. The calibration procedure for the setup involves capturing a standard calibration pattern (e.g., a chessboard) by each camera individually and with the entire sensor system. The camera’s intrinsic (e.g., focal length and distortion coefficients) and extrinsic parameters are required for iMarker pose estimation.
@@
Detection Algorithm. Since the non-CSR regions of the scene lack circular polarization and appear identical in both cameras, it enables the use of simple computer vision techniques (e.g., image subtraction) to detect iMarkers with high contrast, as detailed in [5] and shown below. As shown in Algorithm 1, the iMarker detection procedure requires spatial subtraction of the synchronized frames captured by each camera. Having $\mathbf{F}_{C_{1}}$ and $\mathbf{F}_{C_{2}}$ as the synchronized frame sets captured by calibrated cameras $C_{1}$ and $C_{2}$, each frame $f_{1}\in\mathbf{F}_{C_{1}}$ finds its corresponding synchronized frame $f_{2}\in\mathbf{F}_{C_{2}}$. The next stage is aligning $f_{2}$ based on $f_{1}$ using the alignment parameters $h$ (captured during calibration), where $h$ is the homography matrix generated through matching Oriented FAST and Rotated BRIEF (ORB) features. Subtracting the “aligned” $f_{2}$ from the “original” $f_{1}$ frame results in the final binary image $f_{p}$ containing the input images’ dissimilar parts. Since the CSR region of iMarkers is visible in one frame and blocked in the other, only the patterns created by CSRs are expected to remain in $f_{p}$. Then, a thresholding process with a value of $\theta$, followed by post-processing, is applied to $f_{p}$ to generate a binary frame that enhances the iMarker patterns’ visibility. The final processed frame contains potential iMarkers $m\in\mathbf{M}$, where $\mathbf{M}$ is the set of iMarkers found in $f_{1}$ and $f_{2}$ at time $t$.

### v5

@@ Abstract
Fiducial markers are widely used in robotics for navigation, object recognition, and scene understanding. While offering significant advantages for robots and Augmented Reality (AR) applications, they often disrupt the visual aesthetics of environments, as they are visible to humans, making them unsuitable for many everyday use cases. To address this gap, this paper presents “iMarkers,” innovative, unobtrusive fiducial markers detectable exclusively by robots and AR devices equipped with adequate sensors and detection algorithms. These markers offer high flexibility in production, allowing customization of their visibility range and encoding algorithms to suit various demands. The paper also introduces the hardware designs and open-sourced software algorithms developed for detecting iMarkers, highlighting their adaptability and robustness in the detection and recognition stages. Numerous evaluations have demonstrated the effectiveness of iMarkers relative to conventional (printed) and blended fiducial markers and have confirmed their applicability across diverse robotics scenarios.
@@ 1 Introduction
Fiducial markers, i.e., artificial landmarks with distinguishable graphical patterns, are widely used in robotics and computer vision to offer reliable object identification, localization, and pose estimation [10]. As markers provide better-defined reference points than naturally available ones, they are “simple-yet-effective” solutions when consistent feature matching and precise pose retrieval are essential. Additionally, augmenting information with markers and employing appropriate systems to extract it can benefit various domains, including Augmented Reality (AR) and robotics [5]. For instance, markers can narrow the range of objects a robot needs to recognize and provide real-time pose information for tracking and localization. Thus, decoding data from markers detected using the robot’s camera is less costly than processing the entire visual scene that the robot can see.
Figure 1: A legged robot observing an environment labeled with multiple a) printed fiducial markers and b) the proposed iMarkers, highlighted and magnified.
Although they can convey various advantages, seeing multiple objects labeled with printed markers can be aesthetically undesirable and destructive. In other words, many would prefer to avoid seeing printed tags around them and want the robots to operate effectively without spoiling the zones with these artificial landmarks. Evidence also shows that visible markers can disrupt natural gaze behavior [2]. Consequently, and building upon the idea introduced in [16] for using Cholesteric Spherical Reflectors to create imperceptible landmarks, the authors explored the use of team-developed CSRs and propose a new generation of robotics fiducial markers, dubbed iMarker [^1] https://snt-arg.github.io/iMarkers/. This is achieved by replacing the printing-ink pigment used in common markers with spherical-shell CSRs. Due to the peculiar behavior of CSRs, iMarkers can be indistinguishable to human eyes but observable by robots if a dedicated sensor is provided. Moreover, such markers can be fabricated to be detected only within specific wavelength ranges to avoid visual clutter. Fig. 1 shows that, in qualitative comparisons, iMarkers are hardly noticeable to the naked eye or in standard camera frames, compared with traditional printed fiducial markers.
The motivation of this paper is to demonstrate the role of iMarkers in robotics by presenting their design and detection procedures and benchmarking their applicability in real-world robotic scenarios. Moreover, the work aims to validate the value of iMarkers in simplifying challenging robotic perception tasks, such as detecting transparent and reflective surfaces (e.g., glass, windows, and mirrors). Hence, the leading contributions of the paper are summarized below:
@@
  • Proposing multiple sensor solutions and open-source software algorithms for detection and 6DoF pose estimation of iMarkers of various designs; and
  • Assessing their applicability relative to traditional fiducial markers, with emphasis on scenarios where conventional printed markers are ineffective.
@@ 2.1 Ordinary Fiducial Markers
Fiducial marker libraries are expected to enable rapid detection, reliable recognition, and efficient information decoding. Ordinary markers, mainly printed on paper, can be classified into non-square (designed as circles, point sets, or arbitrary visual patterns), square (or matrix-based), and hybrid variations. In non-square markers, identification and data decoding generally occur w.r.t. the center of their geometric shape. InterSense [14] and WhyCode [13] are circular-shaped libraries in which the correspondence point detection is the tricky part. Dot-shaped libraries, e.g., RUNE-Tag [3], provide fast detection but perform poorly in noisy conditions. Other arbitrary, shape-free markers are primarily used in 6-Degrees-of-Freedom (6DoF) pose estimation. Despite robust detection, they face challenges, including inter-marker-encoded dot confusion and steep viewing-angle recognition.
Square fiducial markers, on the other hand, employ binary codes encoded in matrices and deliver four principal corner points for precise pose estimation. They require lower computational cost for recognition, enabling faster detection algorithms. ARToolkit [11] is a matrix-based marker library that uses template-matching for square marker pattern determination. However, it exhibits high reprojection errors under rotation or camera-angle changes. AprilTag [15] uses gradient-based edge detection and reliable line segment extraction to identify potential tag borders, ensuring accurate marker recognition, even under challenging viewing angles. Its geometrically driven detection pipeline, combined with built-in error-correction mechanisms, significantly reduces false positives caused by background textures and perspective distortions, making AprilTag more robust to rotation, occlusion, and warping than traditional systems. Similarly, Garrido-Jurad et al. introduced ArUco Marker [8], which supports configurable marker dictionary generation with maximized inter-marker Hamming distances. It includes automatic marker detection with error correction via dictionary search, as well as a novel occlusion-handling method using color-based masks for robust pose estimation. The ArUco pipeline uses adaptive thresholding to segment high-contrast regions and separate marker candidates, followed by square-shape isolation and perspective transform computation, which extract bit patterns for predefined dictionary matching.
Despite their widespread use and robust performance, these markers, whether non-square or square, remain visually intrusive and can disrupt the natural aesthetics of environments. This limitation has motivated the development of markers that blend more seamlessly into their surroundings.
@@ 2.2 Blended and Unobtrusive Fiducial Markers
In contrast to paper-based markers, some works propose solutions for scenarios that require specialized detection sensors. In this regard, ArTuga [4] is a multimodal marker for detection in challenging environments through photometric and radiometric sensor fusion. Seedmarker [9] considers the aesthetic properties of markers by embedding them in physical objects, e.g., laser-cut plates and 3D-printed tangibles. ACMarker [19] is another approach that can be detected by custom sensors in underwater robotics applications.
Other approaches focus on integrating less obtrusive markers into environments to minimize their visual impact while maintaining robust detectability. Works in [16, 1] introduced the potential of creating imperceptible visual landmarks by exploiting the optical properties of particular materials. However, they remain primarily focused on materials and optics and do not investigate the applicability of landmarks, such as fiducial markers, for perception robustness, validity, or accuracy. Despite briefly exploring the potential integration of such landmarks in a robotics context in [17], it does not provide a concrete sensing-modality design, a description of detection algorithms, or a systematic experimental evaluation. TRDM [18] uses randomly distributed printed dots on transparent sheets to achieve low visual saliency. Nevertheless, geometric point matching and marker detection are tricky while integrating pairwise relationships among the detected dots. InfraredTags [7] are discreet, durable markers printed and embedded in various objects, providing invisibility and enabling rapid scanning. However, their fabrication is limited to the Infrared (IR) range, reducing design flexibility, and their thin, embedded form makes them difficult to locate visually, complicating alignment of IR illumination and sensing during deployment. BrightMarkers [6] are 3D-printed fluorescent filaments embedded in objects and require Near Infrared (NIR) cameras for detection. Aircode [12] embeds markers directly into the fabrication of objects, enabling seamless integration for use cases such as robotic grasping. Nevertheless, it suffers from a slow decoding time (tens of seconds), which depends on the camera’s viewing angle.
@@ 2.3 Gap Identification
Although the surveyed blended and unobtrusive methodologies have demonstrated effectiveness in robotics and AR tasks by keeping markers discreet, their applicability in real-world scenarios is often constrained by the trade-off between “invisibility” and “detection speed.” In other words, solutions like AirCode achieve high levels of invisibility but suffer from significantly slow detection speeds. Other crucial factors include simplicity, cost-effectiveness, versatility in detection, robust performance, and flexibility in fabrication across various light spectra. These dimensions significantly influence the adoption of such marker solutions compared with straightforward, easily printable traditional markers.
@@
iMarkers offer an innovative, versatile solution to address the gaps mentioned, ensuring transparency, flexibility, and unobtrusiveness. Given appropriate sensors and detection algorithms, they can be fabricated to achieve high contrast for detection while remaining undetectable to the naked eye. Table 1 highlights the characteristics of iMarkers compared to other blended methodologies. Accordingly, although all solutions aim to remain hidden from human perception, iMarkers cover distinct aspects that set them apart:
@@
Core Material: The material used to make iMarkers is CSRs [1], microscopic droplets of Cholesteric Liquid Crystals (CLCs) that self-organize into a helical structure, making them highly effective at reflecting specific wavelengths of light, including IR, Ultraviolet (UV), or visible. Their spherical shape makes them omnidirectional, selective retroreflectors, ensuring detectability regardless of viewing angle. They produce vivid, circularly polarized iridescent colors through selective light reflection, enabling clear separation between CSR and non-CSR regions. As shown in Fig. 2, iMarkers combine CSRs for the marker patterns with transparent UV-curable glue to maintain the structure of the fiducial marker. An alternative fabrication method involves directly spraying CSR shells onto surfaces, a promising approach currently under investigation.
Figure 2: The role of CSRs in the design of iMarkers: (a) an iMarker featuring ArUco patterns filled with CSRs, (b) an iMarker displayed next to visible-range CSR particles, illustrating various shades of green for fabrication.
@@
Fabrication Cost and Efficiency: The cost and simplicity of fabrication significantly impact the widespread adoption of markers. While AirCode requires partitioning and InfraStructs depend on a complex, layered assembly, iMarkers, as shown in Fig. 2, can be fabricated using cost-effective methods such as dispensing or spray coating.
Production Versatility: In contrast with other solutions, iMarkers offer production flexibility, enabling design and detection across the entire light spectrum. As shown in Fig. 3, UV-range iMarkers remain entirely transparent and invisible to the naked eye, even when placed on a transparent surface such as glass. IR-range iMarkers are similarly discreet, though slight scattering or subtle color effects may reveal their presence (to some extent). Even in the visible range, iMarkers can be designed to blend with patterned backgrounds or camouflage against similarly colored surfaces, ensuring minimal visual intrusion.
Sensor Affordability: Detection requirements vary across unobtrusive marker systems: AirCode relies on projector-camera setups, InfraStructs require costly Terahertz scanners, and InfraredTag needs specialized imaging modules with microprocessors. In contrast, iMarkers can be detected using affordable sensors with standard optical components, provided the camera is compatible with the polarization and wavelength characteristics of their CSR material.
@@ 4 Sensor Design and Detection Strategies
Three primary solutions are proposed for iMarker detection, including “dual-vision,” “dynamic single-vision,” and “static single-vision,” further described in the following:
@@ 4.1 Dual-vision Solution
Figure 4: Dual-vision iMarker detection setups using synchronized cameras with circular polarizers and either a cube (a) or plate (b) beamsplitter.
Sensor Design. The first setup is a homogeneous perception system comprising two synchronized, perpendicularly positioned cameras, each facing a different surface of an optical beamsplitter. As shown in Fig. 4, the sensor can be assembled with either a cube or a plate beamsplitter. The beamsplitter equally divides the incoming light (i.e., $50\texttt{:}50$) into two parts: the transmitted and the reflected light. By mounting the cameras and the beamsplitter in a cage, one camera captures the scene using the transmitted light, while the other captures it using the reflected light, making two perspectives with minimal variation. Attaching identical circular polarizers to each camera’s lens effectively blocks light with the opposite circular polarization, ensuring only the desired handedness is detected. Due to the inversion of circular polarization upon reflection at the beamsplitter, the camera capturing the reflected light detects the “opposite” circular polarization to that of the camera capturing the transmitted light. Hence, the setup simultaneously captures the same scene with two cameras, ensuring that the CSR regions of iMarkers in the sensor’s Field of View (FoV) appear blocked in one camera while remaining visible in the other (due to the opposing circular polarization of the CSR reflections).
Note that the cameras in this setup must be synchronized to capture the same scene simultaneously (frame $f_{t}^{C_{1}}$ captured by camera $C_{1}$ at time $t$ should depict the same scene as frame $f_{t}^{C_{2}}$ of camera $C_{2}$). As the cameras’ frame rate and shutter speed are vital for high-speed robotic applications (e.g., drone navigation), integrating global-shutter cameras in this sensor is highly recommended. The calibration procedure involves capturing a standard calibration pattern (e.g., a chessboard) by each camera. The camera’s intrinsic (e.g., focal length and distortion coefficients) and extrinsic parameters are required for iMarker pose estimation.
@@
Detection Algorithm. Since the non-CSR regions of the scene lack circular polarization and appear identical in both cameras, simple frame subtraction can detect iMarkers with high contrast, as shown in Algorithm 1. Accordingly, the iMarker detection procedure requires spatial subtraction of the synchronized frames captured by each camera. Having $\mathbf{F}_{C_{1}}$ and $\mathbf{F}_{C_{2}}$ as the synchronized frame sets captured by calibrated cameras $C_{1}$ and $C_{2}$, each frame $f_{1}\in\mathbf{F}_{C_{1}}$ finds its corresponding synchronized frame $f_{2}\in\mathbf{F}_{C_{2}}$. The next stage is to align $f_{2}$ with respect to $f_{1}$ using the alignment parameters $h$ (captured during calibration), where $h$ is the homography matrix computed from matching Oriented FAST and Rotated BRIEF (ORB) features. Subtracting the “aligned” $f_{2}$ from the “original” $f_{1}$ frame results in the final binary image $f_{p}$ containing the input images’ dissimilar parts. Since the CSR region of iMarkers is visible in one frame and blocked in the other, only the patterns created by CSRs are expected to remain in $f_{p}$. Then, a thresholding process with a value of $\theta$, followed by post-processing, is applied to $f_{p}$ to generate a binary frame that enhances the iMarker patterns’ visibility. The final processed frame contains potential iMarkers $\mathbf{M}$ at time $t$.